OpenStudy (anonymous):

Simplify the expression 4^-2 ^6 ------- 4^2 2^2 answers A. 64 B. 1/16 C.1 D.1/2

OpenStudy (neer2890):

is it\[\frac{ 4^{-2^{6}} }{ 4^{2}*2^{2} }\]??

OpenStudy (anonymous):

Idk ;/

OpenStudy (neer2890):

then what's your question..?

OpenStudy (neer2890):

i assume this is your question\[\frac{ 4^{-2}*2^{6} }{ 4^{2}*2^{2} }\] =\[\frac{ 2^{6} }{ 4^{2}*4^{2}*2^{2} }=\frac{ 2^{4} }{ 4^{4} }=\frac{ 1 }{ 16 }\]

OpenStudy (anonymous):

@neer2890 how did you make the exponents appear over the number ?

OpenStudy (neer2890):

use equation button..!

OpenStudy (anonymous):

Oh thanks!!

OpenStudy (neer2890):

you're welcome...:)

OpenStudy (anonymous):

ooh yeaa thanks